# Long distance propagation of light in random media with partially coherent sources

### v1

@@ 1 Introduction
However the presence of random sources leads to more complex mixing regimes with scintillation index different from unity [5, 16]. It has been observed that the superior performance of such sources towards scintillation reduction is in part due to temporal averaging at the detector. If the averaging time of the detector is much smaller than the decorrelation time of the source, scintillation can potentially be higher for partially coherent sources due to the additional randomness of the source [5, 18, 16]. The goal of this paper is to provide a description of the limiting distribution of the wave field due to partially coherent sources under the kinetic and diffusive scalings of the scintillation regime. We will also use this to characterize the long distance statistics of time averaged intensity in terms of the coherence length scales of the source and time averaging at the detector.
@@
$\Big{(}\partial^{2}_{z}+\Delta_{x}+k_{0}^{2}n^{2}(z,x)\Big{)}p(z,x)=-u_{0}(x)\delta_{0}(z)\,,$
@@
where $t$ reflects the time scale of the random fluctuations in the incident beam. We assume that this time scale is much larger than the time it takes an incident beam to reach a receiver at distance $z>0$. Thus, $t$ can be assumed to be a frozen parameter in (1). The randomness in the medium is encoded in $B(z,x;t)$, a Gaussian process over $[0,\infty)\times\mathbb{R}^{d}\times[0,\infty)$ with statistics
@@
with $F(-t)=F(t)\in{\mathsf{L}}^{1}(\mathbb{R})\cap{\mathsf{L}}^{\infty}(\mathbb{R})$ and $10\leq F(t)\leq F(0)=$ 1 and $0<\tau_{s}\leq 1$ models the time scale of the random source. To simplify notation, we will simply replace the above term by $F(\tau-\tau^{\prime})$. We will observe that only for $\tau_{s}$ sufficiently small compared to the detector time scale ${\mathsf{T}}$ do we expect a small scintillation index.
@@
Here, $\Gamma(r,-\sigma)=\Gamma(r,\sigma)\in\mathcal{S}(\mathbb{R}^{2d})$ and $0<\theta\leq 1$ a parameter (independent of $\varepsilon$) that controls the spatial correlation length of the source with respect to the overall width of the beam. As in [4], the statistics of the wavefield satisfy an asymptotic limit in the kinetic and diffusive regimes as $\varepsilon\to 0$ provided that the incident beam is sufficiently broad. This is modeled by a choice of parameter $\beta\geq 1$ with the incident field a broad beam given as a (smooth) function of $\varepsilon^{\beta}x$ and hence a diameter of order $\varepsilon^{-\beta}$. The case $\beta=1$ generates a richer limiting diffusive model. The case $\beta>1$ models very broad incident beams as a good approximation to plane waves.
@@
[…]
@@
$\mathbb{E}[\prod\limits_{j=1}^{p}u_{0}^{\varepsilon}(x_{j};t_{j})\prod\limits_{l=1}^{q}u_{0}^{\varepsilon*}(y_{l};t^{\prime}_{l})]=\begin{cases}0,\quad p\neq q\\ \sum\limits_{\pi_{p}}\prod\limits_{j=1}^{p}F(t_{j}-t^{\prime}_{\pi_{p}(j)})J^{\varepsilon}(x_{j},y_{\pi_{p}(j)}),\quad p=q\,,\end{cases}\end{aligned}$ (7)
@@
where $N^{\varepsilon}_{p,q}(z)$ and $E^{\varepsilon}_{p,q}(z)$ are operators acting on $\mathcal{M}_{B}(\mathbb{R}^{(p+q)d})$ and $\Pi^{\varepsilon}_{p,q}(z,v)$ is a unit magnitude complex exponential defined explicitly below in (44). $N^{\varepsilon}_{p,q}(z)$ (defined in (51)) is an operator constructed out of first and second moment type solution operators while the remainder satisfies $\|E_{p,q}^{\varepsilon}(z)\|\leq c(p,q,z)\varepsilon^{\frac{1}{3}}$.
@@ 3 Main results
$\mathbb{E}[I(z,r)^{p}]=\begin{cases}\Gamma(r,0)^{p}p!^{2},\quad&\beta>1\\ p![G_{p}(z^{3},\cdot)\ast S_{p}^{I}(\cdot)](r,\ldots,r),\quad&{\beta=1}\\ p!\mathbb{E}[I](z,r)^{p},\quad&\beta=1,\ \theta\to 0\,,\end{cases}$ (22)
@@
The proofs of the above theorems and corollaries are presented in Sections 6 and 7 while the tightness result is presented in an Appendix. These proofs follow from a careful analysis of the limiting moments of the random vector $\Phi^{\varepsilon}$. The $p+p$th moments of such random vectors do not grow faster than $C^{p}p!^{2}$, so that the limiting distributions are characterized by their moments [6].
Additional smoothness assumptions on the source and medium allow to show the tightness criterion (30) and obtain that the process $\phi^{\varepsilon}(z,r)$ in fact converges in distribution as probability measures on $C^{0}(\mathbb{R}^{d}\times\mathbb{R})$ [6, 23]. These results generalize those obtained in [4] for coherent sources to the setting of partially coherent incident beams.
@@ 4 Interpretation and illustrations of the results
In the kinetic regime, $\mathbb{E}[Z_{2}](z,x)=e^{-\frac{k_{0}^{2}R(0)z}{8}}$ and $Z(z,x)$ is a mean zero complex Gaussian with $\mathbb{E}[Z(z,x)Z^{\ast}(z,y)]=\mathcal{R}(y-x,0)e^{-\frac{k_{0}^{2}R(0)z}{4}}$ and $\mathcal{R}$ defined in (46).
@@
Scintillations computed in these examples also align with the limiting expressions in (28). The time averaged scintillation for various configurations of $\tau_{s}/{\mathsf{T}}$ and $\theta$ are plotted in Fig 1 and Fig 2. As predicted from the calculations, increasing time averaging $\tau_{s}/{\mathsf{T}}\to 0$ and reducing coherence length $\theta\to 0$ both clearly favour a reduction in scintillation. The rest of the paper is devoted to the proof of our main results. We recall that in order to simplify notation, we systematically replaced $F(\frac{t}{\tau_{s}})$ by $F(t)$.
@@ 5 Two-point statistics of the wave field
This can be solved explicitly [14, 4] to give
$m_{1,1}^{\varepsilon}(z,r,x,y;t,t^{\prime})$ $=F(t-t^{\prime})\int\limits_{\mathbb{R}^{2d}}\hat{\Gamma}^{\varepsilon}(\zeta,\xi)e^{i\varepsilon^{-\beta}r\cdot\zeta}e^{i\eta\xi\cdot(x-y)}e^{i\eta\zeta\cdot\frac{(x+y)}{2}}e^{-\frac{iz\eta}{k_{0}\varepsilon}\xi\cdot\zeta}\mathcal{Q}\big{(}\eta(y-x),\frac{\eta\zeta}{\varepsilon}\big{)}\frac{\mathrm{d}\xi\mathrm{d}\zeta}{(2\pi)^{2d}}\,,$
@@
$\mathcal{Q}(\tau,\tau^{\prime})=e^{\frac{k_{0}^{2}z}{4\eta^{2}}\int\limits_{0}^{1}Q\big{(}\tau+\frac{\tau^{\prime}sz}{k_{0}}\big{)}\mathrm{d}s},\quad Q(x)=R(x)-R(0)\,.$ (32)
@@
$m_{1,1}^{\varepsilon}(z,r,x,y;t,t^{\prime})$ $=F(t-t^{\prime})\!\!\int\limits_{\mathbb{R}^{2d}}\!\!\hat{\Gamma}(\zeta,\xi)e^{ir\cdot\zeta}e^{i\eta\varepsilon^{\beta}\xi\cdot\frac{(x-y)}{\theta}}e^{i\eta\varepsilon^{\beta}\zeta\cdot\frac{(x+y)}{2}}e^{-\frac{iz\eta\varepsilon^{2\beta-1}}{k_{0}\theta}\xi\cdot\zeta}\mathcal{Q}\big{(}\eta(y-x),\eta\varepsilon^{\beta-1}\zeta\big{)}\frac{\mathrm{d}\xi\mathrm{d}\zeta}{(2\pi)^{2d}}.$
@@
$\lim_{\varepsilon\to 0}\mathcal{Q}\big{(}\eta(y-x),\eta\varepsilon^{\beta-1}\zeta\big{)}=\begin{cases}\mathcal{Q}(y-x,0),\quad\beta>1\\ \mathcal{Q}(y-x,\zeta),\quad\beta=1\,.\end{cases}$ (33)
@@
$M_{1,1}(z,r,x,y;t,t^{\prime})=\begin{cases}F(t-t^{\prime})\Gamma(r,0)\mathcal{Q}(y-x,0),\quad\beta>1\\ F(t-t^{\prime})\int\limits_{\mathbb{R}^{2d}}\Gamma(r^{\prime},0)\mathcal{Q}(y-x,\zeta)e^{i\zeta\cdot(r-r^{\prime})}\frac{\mathrm{d}\zeta\mathrm{d}r^{\prime}}{(2\pi)^{d}},\quad\beta=1\,.\end{cases}$ (34)
@@
$\lim_{\varepsilon\to 0}\mathcal{Q}\big{(}\eta(y-x),\eta\varepsilon^{\beta-1}\zeta\big{)}=\begin{cases}\exp\Big{(}\frac{k_{0}^{2}z}{8}(y-x)^{\mathrm{t}}\Xi(y-x)\Big{)},\quad\beta>1\\ \exp\Big{(}\frac{k_{0}^{2}z}{8}\int\limits_{0}^{1}\big{(}y-x+\frac{\zeta sz}{k_{0}}\big{)}^{\mathrm{t}}\Xi\big{(}y-x+\frac{\zeta sz}{k_{0}}\big{)}\mathrm{d}s\Big{)},\quad\beta=1\,,\end{cases}$ (35)
@@ 6 Higher moments of the wave field
$\mu^{\varepsilon}_{p,q}(z=0,X,Y;T)=\begin{cases}0,\quad p\neq q\\ \sum\limits_{\pi_{p}}\prod\limits_{j=1}^{p}F(t_{j}-t^{\prime}_{\pi_{p}(j)})J^{\varepsilon}(x_{j},y_{\pi_{p}(j)}),\quad p=q\,.\end{cases}\end{aligned}$ (41)
@@
$\hat{\mu}^{\varepsilon}_{p,q}(z=0,v;T)=\begin{cases}0,\quad p\neq q\\ \sum\limits_{\pi_{p}}\prod\limits_{j=1}^{p}F(t_{j}-t^{\prime}_{\pi_{p}(j)})\hat{J}^{\varepsilon}(\xi_{j},\zeta_{\pi_{p}(j)}),\quad p=q\,.\end{cases}\end{aligned}$ (43)
@@
$\mathcal{R}(\tau,\tau^{\prime})=\exp\Big{(}{\frac{k_{0}^{2}z}{4\eta^{2}}\int\limits_{0}^{1}R\big{(}\tau+\frac{\tau^{\prime}sz}{k_{0}}\big{)}\mathrm{d}s}\Big{)}.$ (46)
@@
$U^{\varepsilon}_{j,l}\rho(z,v)=\int\limits_{\mathbb{R}^{2d}}e^{-iw\cdot k}\mathcal{R}\big{(}w,\frac{\eta}{\varepsilon}A_{j,l}^{\mathrm{t}}\Theta v\big{)}\rho(0,v-A_{j,l}k)\frac{\mathrm{d}k\mathrm{d}w}{(2\pi)^{d}}\,.$ (48)
@@
$\rho^{\prime}(z,\xi_{j}^{\prime},\zeta_{l}^{\prime})$ $=\int\limits_{\mathbb{R}^{2d}}e^{-iw\cdot(\xi^{\prime}_{j}-k)}\rho^{\prime}(0,k,\zeta^{\prime}_{l})\mathcal{R}\Big{(}w,\frac{\eta z\zeta^{\prime}_{l}}{\varepsilon k_{0}}\Big{)}\frac{\mathrm{d}k\mathrm{d}w}{(2\pi)^{d}}\,.$
@@
$\rho(z,\xi_{j},\zeta_{l})$ $=\int\limits_{\mathbb{R}^{2d}}e^{-iw\cdot\big{(}\frac{\xi_{j}+\zeta_{l}}{2}-k\big{)}}\rho_{0}\Big{(}k+\frac{\xi_{j}-\zeta_{l}}{2},k-\frac{\xi_{j}-\zeta_{l}}{2}\Big{)}\mathcal{R}\Big{(}w,\frac{\eta z(\xi_{j}-\zeta_{l})}{\varepsilon k_{0}}\Big{)}\frac{\mathrm{d}k\mathrm{d}w}{(2\pi)^{d}}\,.$
@@
The details can be found in [4, Theorem 4.2]. Since $T$ is a fixed vector of parameters that do not affect (42) and since $\hat{\mu}_{p,q}^{\varepsilon}(0,v;T)$ is integrable, Lemma 6.2 allows one to decompose $\hat{\mu}_{p,q}^{\varepsilon}$ as
@@
Due to the form of the source (43), $\hat{\mu}_{p,q}^{\varepsilon}$ is equal to zero trivially when $p\neq q$. For $p=q$, mapping this back to the physical domain gives a slight reformulation of [4, Theorem 4.12].
@@
$\tilde{N}_{p,p}^{\varepsilon}(z,X,Y;T)$ $=U_{\eta}^{p}\sum_{\kappa=1}^{K}\int\limits_{\mathbb{R}^{2pd}}e^{iv^{\mathrm{t}}\Theta{\mathsf{x}}}e^{-\frac{iz\eta}{2k_{0}\varepsilon}v^{\mathrm{t}}\Theta v}\prod\limits_{\gamma\in\Lambda_{\kappa}}\Big{(}\mathcal{R}(-A_{\gamma}^{\mathrm{t}}\Theta{\mathsf{x}},\frac{\eta}{\varepsilon}A_{\gamma}^{\mathrm{t}}\Theta v)-1\Big{)}\hat{\mu}^{\varepsilon}_{p,p}(0,v;T)\prod_{j=1}^{p}\frac{\mathrm{d}\xi_{j}\mathrm{d}\zeta_{j}}{(2\pi)^{2d}}$ (55)
@@
$\times$ $\prod_{\gamma\in\Lambda_{\kappa}}\big{(}\mathcal{R}(w_{\gamma_{1}},\frac{\eta}{\varepsilon}A_{\gamma}^{\mathrm{t}}\Theta v)-1\big{)}\frac{\mathrm{d}w_{\gamma_{1}}\mathrm{d}k_{\gamma_{2}}}{(2\pi)^{d}}+\hat{\mu}^{\varepsilon}_{p,p}(0,v;T)\Big{)}\prod_{j=1}^{p}\frac{\mathrm{d}\xi_{j}\mathrm{d}\zeta_{j}}{(2\pi)^{2d}}\,.$
@@
×$\qquad\times\prod\limits_{\gamma\in\Lambda_{\kappa}}\Big{(}\mathcal{R}(w_{\gamma_{1}},\frac{\eta}{\varepsilon}A_{\gamma}^{\mathrm{t}}\Theta v)-1\Big{)}\hat{\mu}^{\varepsilon}_{p,p}(0,v;T)\prod_{\gamma\in\Lambda_{\kappa}}\frac{\mathrm{d}w_{\gamma_{1}}\mathrm{d}k_{\gamma_{2}}}{(2\pi)^{d}}\prod_{j=1}^{p}\frac{\mathrm{d}\xi_{j}\mathrm{d}\zeta_{j}}{(2\pi)^{2d}}$
@@
We obtain the same bound in the physical domain for $\|\mu^{\varepsilon}_{p,p}(z)\|_{\infty}$ uniformly in $(X,Y)$ and $T$, which means we can uniquely characterize limiting distributions through their moments, provided the moments converge [6].
@@
Here $\Lambda_{\kappa,1}$ denotes the set of all indices in the first position in the pairs in $\Lambda_{\kappa}$ and $\Lambda_{\kappa,2}$ denotes the set of all indices in the second position.
@@
×$\times\prod\limits_{\gamma\in\Lambda_{\kappa}}\Big{(}\mathcal{R}\big{(}\eta(y_{\gamma_{2}}-x_{\gamma_{1}}),\eta\varepsilon^{\beta-1}(\xi_{\gamma_{1}}-\zeta_{\gamma_{2}})\big{)}-1\Big{)}\hat{S_{p}}(v;T)\prod_{j=1}^{p}\frac{\mathrm{d}\xi_{j}\mathrm{d}\zeta_{j}}{(2\pi)^{2d}}$

### v2

@@ 1 Introduction
However the presence of random sources leads to more complex mixing regimes with scintillation index different from unity [1, 7, 8, 15, 21]. It has been observed that the superior performance of such sources towards scintillation reduction is in part due to temporal averaging at the detector. If the averaging time of the detector is much smaller than the decorrelation time of the source, scintillation can potentially be higher for partially coherent sources due to the additional randomness of the source [8, 23, 21]. The goal of this paper is to provide a description of the limiting distribution of the wave field due to partially coherent sources under the kinetic and diffusive scalings of the scintillation regime. We will also use this to characterize the long distance statistics of time averaged intensity in terms of the coherence length scales of the source and time averaging at the detector.
@@
$\Big{(}\partial^{2}_{z}+\Delta_{x}+k_{0}^{2}n^{2}(z,x)\Big{)}p(z,x)={\color[rgb]{0,0,0}+}u_{0}(x){\color[rgb]{0,0,0}\delta^{\prime}_{0}(z)}\,$,
@@
where $t$ reflects the time scale of the random fluctuations in the incident beam. We assume that this time scale is much larger than the time it takes an incident beam to reach a receiver at distance $z>0$. Thus, $t$ can be assumed to be a frozen parameter in (1). The randomness in the medium is encoded in $B(z,x;t)$, a Gaussian process over $0,\infty)\times\mathbb{R}^{d}\times[0,\infty)$ with statistics
@@
with $F(-t)=F(t)\in{\mathsf{L}}^{1}(\mathbb{R})\cap{\mathsf{L}}^{\infty}(\mathbb{R})$ and for simplicity, we assume $0\leq|F(t)|\leq F(0)=1$ and $0<\tau_{s}\leq 1$ models the time scale of the random source. To simplify notation, we will simply replace the above term by $F(t-t^{\prime})$. We will observe that only for $\tau_{s}$ sufficiently small compared to the detector time scale $\mathsf{T}$} do we expect a small scintillation index.
@@
We assume that $\Gamma$ depends only on the barrycentre $r=\frac{x+y}{2}$ and the magnitude of separation between the points $\sigma=x-y$ so that the covariance $J(x,y)=\Gamma\big{(}\frac{x+y}{2},x-y\big{)}=\Gamma\big{(}\frac{x+y}{2},|x-y|\big{)}$. Here, $\Gamma(r,-\sigma)=\Gamma(r,\sigma)\in\mathcal{S}(\mathbb{R}^{2d})$ and $0<\theta\leq 1$ a parameter (independent of $\varepsilon$) that controls the spatial correlation length of the source with respect to the overall width of the beam. As in [5], the statistics of the wavefield satisfy an asymptotic limit in the kinetic and diffusive regimes as $\varepsilon\to 0$ provided that the incident beam is sufficiently broad. This is modeled by a choice of parameter $\beta\geq 1$ with the incident field a broad beam given as a (smooth) function of $\varepsilon^{\beta}x$ and hence a diameter of order $\varepsilon^{-\beta}$. The case $\beta=1$ generates a richer limiting diffusive model. The case $\beta>1$ models very broad incident beams as a good approximation to plane waves.
@@
We denote by $\mathcal{M}_{B}(\mathbb{R}^{nd})$, the Banach space of finite signed measures on $\mathbb{R}^{nd}$ for $n$ an integer, equipped with the total variation (TV) norm $\|\cdot\|$; for measures with a density with respect to the Lebesgue measure, then that norm is the ${\mathsf{L}}^{1}(\mathbb{R}^{nd})$ norm of that density. Throughout the paper, $\|\cdot\|$ is used for the TV norm of bounded Radon measures or for the operator norm of operators acting on such measures.
@@
[…]
@@
where $v=(\xi_{1},\cdots,\xi_{p},\zeta_{1},\cdots,\zeta_{q})$ denotes the vector of Fourier variables dual to $(X,Y)$, $N^{\varepsilon}_{p,q}(z)$ and $E^{\varepsilon}_{p,q}(z)$ are operators acting on $\mathcal{M}_{B}(\mathbb{R}^{(p+q)d})$ and $\Pi^{\varepsilon}_{p,q}(z,v)$ is a unit magnitude complex exponential defined explicitly below in (44). $N^{\varepsilon}_{p,q}(z)$ (defined in (51)) is an operator constructed out of first and second moment type solution operators while the remainder satisfies $\|E_{p,q}^{\varepsilon}(z)\|\leq c(p,q,z)\varepsilon^{\frac{1}{3}}$.
@@ 3 Main results
[…]
@@
The proofs of the above theorems and corollaries are presented in Sections 6 and 7 while the tightness result is presented in an Appendix. These proofs follow from a careful analysis of the limiting moments of the random vector $\Phi^{\varepsilon}$. The $p+p$th moments of such random vectors do not grow faster than $C^{p}p!^{2}$, so that the limiting distributions are characterized by their moments [9, 14].
Additional smoothness assumptions on the source and medium allow to show the tightness criterion (30) and obtain that for $(z,r)$ fixed, the process $\phi^{\varepsilon}(z,r)$ in fact converges in distribution as probability measures on $C^{0}(\mathbb{R}^{d}\times\mathbb{R})$ [9, 28]. The finite dimensional distributions are independent at different values of $r$. This may be verified as in [5, Corollary 2.6] since it morally corresponds to an infinite distance compared to the effective correlation length $\eta$. These results generalize those obtained in [5] for coherent sources to the setting of partially coherent incident beams.
@@ 4 Interpretation and illustrations of the results
In the kinetic regime, $\mathbb{E}[Z_{2}](z,x)=e^{-\frac{k_{0}^{2}R(0)z}{8}}$ and $Z(z,x)$ is a mean zero complex Gaussian with $\mathbb{E}[Z(z,x)Z^{\ast}(z,y)]={\color[rgb]{0,0,0}\mathcal{R}_{z}}(y-x,0)e^{-\frac{k_{0}^{2}R(0)z}{4}}$ and $\color[rgb]{0,0,0}\mathcal{R}_{z}$} defined in (46).
@@
Scintillations computed in these examples also align with the limiting expressions in (28). The time averaged scintillation for various configurations of $\tau_{s}/{\mathsf{T}}$ and $\theta$ are plotted in Fig 1 and Fig 2. As predicted from the calculations, increasing time averaging $\tau_{s}/{\mathsf{T}}\to 0$ and reducing coherence length $\theta\to 0$ both clearly favour a reduction in scintillation.The rest of the paper is devoted to the proof of our main results.
@@ 5 Two-point statistics of the wave field
We recall that in order to simplify notation, we systematically replaced $F(\frac{t}{\tau_{s}})$ by $F(t)$. The equations for $\mu^{\varepsilon}_{1,1}$ and $m^{\varepsilon}_{1,1}$ can be solved explicitly [19, 5] to give
$m_{1,1}^{\varepsilon}(z,r,x,y;t,t^{\prime})$ $=F(t-t^{\prime})\int\limits_{\mathbb{R}^{2d}}\hat{\Gamma}^{\varepsilon}(\zeta,\xi)e^{i\varepsilon^{-\beta}r\cdot\zeta}e^{i\eta\xi\cdot(x-y)}e^{i\eta\zeta\cdot\frac{(x+y)}{2}}e^{-\frac{iz\eta}{k_{0}\varepsilon}\xi\cdot\zeta}{\color[rgb]{0,0,0}\mathcal{Q}_{z}}\big{(}\eta(y-x),\frac{\eta\zeta}{\varepsilon}\big{)}\frac{\mathrm{d}\xi\mathrm{d}\zeta}{(2\pi)^{2d}}\,$,
@@
${\color[rgb]{0,0,0}\mathcal{Q}_{z}}(\tau,\tau^{\prime})=e^{\frac{k_{0}^{2}z}{4\eta^{2}}\int\limits_{0}^{1}Q\big{(}\tau+\frac{\tau^{\prime}sz}{k_{0}}\big{)}\mathrm{d}s},\quad Q(x)=R(x)-R(0)\,.$ (32)
@@
$m_{1,1}^{\varepsilon}(z,r,x,y;t,t^{\prime})$ $=F(t-t^{\prime})\!\!\int\limits_{\mathbb{R}^{2d}}\!\!\hat{\Gamma}(\zeta,\xi)e^{ir\cdot\zeta}e^{i\eta\varepsilon^{\beta}\xi\cdot\frac{(x-y)}{\theta}}e^{i\eta\varepsilon^{\beta}\zeta\cdot\frac{(x+y)}{2}}e^{-\frac{iz\eta\varepsilon^{2\beta-1}}{k_{0}\theta}\xi\cdot\zeta}{\color[rgb]{0,0,0}\mathcal{Q}_{z}}\big{(}\eta(y-x),\eta\varepsilon^{\beta-1}\zeta\big{)}\frac{\mathrm{d}\xi\mathrm{d}\zeta}{(2\pi)^{2d}}.$
@@
$\lim_{\varepsilon\to 0}{\color[rgb]{0,0,0}\mathcal{Q}_{z}}\big{(}\eta(y-x),\eta\varepsilon^{\beta-1}\zeta\big{)}=\begin{cases}{\color[rgb]{0,0,0}\mathcal{Q}_{z}}(y-x,0),\quad\beta>1\\ {\color[rgb]{0,0,0}\mathcal{Q}_{z}}(y-x,\zeta),\quad\beta=1\,.\end{cases}$ (33)
@@
$M_{1,1}(z,r,x,y;t,t^{\prime})=\begin{cases}F(t-t^{\prime})\Gamma(r,0){\color[rgb]{0,0,0}\mathcal{Q}_{z}}(y-x,0),\quad\beta>1\\ F(t-t^{\prime})\int\limits_{\mathbb{R}^{2d}}\Gamma(r^{\prime},0){\color[rgb]{0,0,0}\mathcal{Q}_{z}}(y-x,\zeta)e^{i\zeta\cdot(r-r^{\prime})}\frac{\mathrm{d}\zeta\mathrm{d}r^{\prime}}{(2\pi)^{d}},\quad\beta=1\,.\end{cases}$ (34)
@@
[…]
@@ 6 Higher moments of the wave field
$\begin{aligned} \mu^{\varepsilon}_{p,q}(z=0,X,Y;T)=\begin{cases}0,\quad p\neq q\\ \sum\limits_{\pi_{p}}\prod\limits_{j=1}^{p}F(t_{j}-t^{\prime}_{\pi_{p}(j)})J^{\varepsilon}(x_{j},y_{\pi_{p}(j)}),\quad p=q\,.\end{cases}\end{aligned}$ (41)
@@
$\begin{aligned} \hat{\mu}^{\varepsilon}_{p,q}(z=0,v;T)=\begin{cases}0,\quad p\neq q\\ \sum\limits_{\pi_{p}}\prod\limits_{j=1}^{p}F(t_{j}-t^{\prime}_{\pi_{p}(j)})\hat{J}^{\varepsilon}(\xi_{j},\zeta_{\pi_{p}(j)}),\quad p=q\,.\end{cases}\end{aligned}$ (43)
@@
${\color[rgb]{0,0,0}\mathcal{R}_{z}}(\tau,\tau^{\prime})=\exp\Big{(}{\frac{k_{0}^{2}z}{4\eta^{2}}\int\limits_{0}^{1}R\big{(}\tau+\frac{\tau^{\prime}sz}{k_{0}}\big{)}\mathrm{d}s}\Big{)}.$ (46)
@@
$U^{\varepsilon}_{j,l}\rho(z,v)=\int\limits_{\mathbb{R}^{2d}}e^{-iw\cdot k}{\color[rgb]{0,0,0}\mathcal{R}_{z}}\big{(}w,\frac{\eta}{\varepsilon}A_{j,l}^{\mathrm{t}}\Theta v\big{)}\rho(0,v-A_{j,l}k)\frac{\mathrm{d}k\mathrm{d}w}{(2\pi)^{d}}\,.$ (48)
@@
$\rho^{\prime}(z,\xi_{j}^{\prime},\zeta_{l}^{\prime})$ $=\int\limits_{\mathbb{R}^{2d}}e^{-iw\cdot(\xi^{\prime}_{j}-k)}\rho^{\prime}(0,k,\zeta^{\prime}_{l}){\color[rgb]{0,0,0}\mathcal{R}_{z}}\Big{(}w,\frac{\eta z\zeta^{\prime}_{l}}{\varepsilon k_{0}}\Big{)}\frac{\mathrm{d}k\mathrm{d}w}{(2\pi)^{d}}\,.$
@@
$\rho(z,\xi_{j},\zeta_{l})$ $=\int\limits_{\mathbb{R}^{2d}}e^{-iw\cdot\big{(}\frac{\xi_{j}+\zeta_{l}}{2}-k\big{)}}\rho_{0}\Big{(}k+\frac{\xi_{j}-\zeta_{l}}{2},k-\frac{\xi_{j}-\zeta_{l}}{2}\Big{)}{\color[rgb]{0,0,0}\mathcal{R}_{z}}\Big{(}w,\frac{\eta z(\xi_{j}-\zeta_{l})}{\varepsilon k_{0}}\Big{)}\frac{\mathrm{d}k\mathrm{d}w}{(2\pi)^{d}}\,.$
@@
The details can be found in [5, Theorem 4.2]. Since $T$ is a fixed vector of parameters that do not affect (42) and since $\hat{\mu}_{p,q}^{\varepsilon}(0,v;T)$ is integrable, Lemma 6.2 allows one to decompose $\hat{\mu}_{p,q}^{\varepsilon}$ as
@@
Due to the form of the source (43), $\hat{\mu}_{p,q}^{\varepsilon}$ is equal to zero trivially when $p\neq q$. For $p=q$, mapping this back to the physical domain gives a slight reformulation of [5, Theorem 4.12].
@@
$\tilde{N}_{p,p}^{\varepsilon}(z,X,Y;T)$ $=U_{\eta}^{p}\sum_{\kappa=1}^{K}\int\limits_{\mathbb{R}^{2pd}}e^{iv^{\mathrm{t}}\Theta{\mathsf{x}}}e^{-\frac{iz\eta}{2k_{0}\varepsilon}v^{\mathrm{t}}\Theta v}\prod\limits_{\gamma\in\Lambda_{\kappa}}\Big{(}{\color[rgb]{0,0,0}\mathcal{R}_{z}}(-A_{\gamma}^{\mathrm{t}}\Theta{\mathsf{x}},\frac{\eta}{\varepsilon}A_{\gamma}^{\mathrm{t}}\Theta v)-1\Big{)}\hat{\mu}^{\varepsilon}_{p,p}(0,v;T)\prod_{j=1}^{p}\frac{\mathrm{d}\xi_{j}\mathrm{d}\zeta_{j}}{(2\pi)^{2d}}$ (55)
@@
$\times$ $\prod_{\gamma\in\Lambda_{\kappa}}\big{(}{\color[rgb]{0,0,0}\mathcal{R}_{z}}(w_{\gamma_{1}},\frac{\eta}{\varepsilon}A_{\gamma}^{\mathrm{t}}\Theta v)-1\big{)}\frac{\mathrm{d}w_{\gamma_{1}}\mathrm{d}k_{\gamma_{2}}}{(2\pi)^{d}}+\hat{\mu}^{\varepsilon}_{p,p}(0,v;T)\Big{)}\prod_{j=1}^{p}\frac{\mathrm{d}\xi_{j}\mathrm{d}\zeta_{j}}{(2\pi)^{2d}}\,.$
@@
$\qquad\times\prod\limits_{\gamma\in\Lambda_{\kappa}}\Big{(}{\color[rgb]{0,0,0}\mathcal{R}_{z}}(w_{\gamma_{1}},\frac{\eta}{\varepsilon}A_{\gamma}^{\mathrm{t}}\Theta v)-1\Big{)}\hat{\mu}^{\varepsilon}_{p,p}(0,v;T)\prod_{\gamma\in\Lambda_{\kappa}}\frac{\mathrm{d}w_{\gamma_{1}}\mathrm{d}k_{\gamma_{2}}}{(2\pi)^{d}}\prod_{j=1}^{p}\frac{\mathrm{d}\xi_{j}\mathrm{d}\zeta_{j}}{(2\pi)^{2d}}$
@@
We obtain the same bound in the physical domain for $\|\mu^{\varepsilon}_{p,p}(z)\|_{\infty}$ uniformly in $(X,Y)$ and $T$, which means we can uniquely characterize limiting distributions through their moments, provided the moments converge [9, 14].
@@
Here $\Lambda_{\kappa,1}$ denotes the set of all indices in the first position in the pairs in $\Lambda_{\kappa}$ and $\Lambda_{\kappa,2}$ denotes the set of all indices in the second position. We note that the functional $\mathcal{F}$ is a relationship between raw moments. This is consistent with the Gaussian summation rule that centred higher order even moments can be written in terms of products of centred second moments.
@@
$\times\prod\limits_{\gamma\in\Lambda_{\kappa}}\Big{(}{\color[rgb]{0,0,0}\mathcal{R}_{z}}\big{(}\eta(y_{\gamma_{2}}-x_{\gamma_{1}}),\eta\varepsilon^{\beta-1}(\xi_{\gamma_{1}}-\zeta_{\gamma_{2}})\big{)}-1\Big{)}\hat{S_{p}}(v;T)\prod_{j=1}^{p}\frac{\mathrm{d}\xi_{j}\mathrm{d}\zeta_{j}}{(2\pi)^{2d}}$